katkit25:

can you solve (4*20) + ( 1/2(4*5) )

pooja195:

Hi Welcome to questioncove! Use order of operations

pooja195:

PEMDAS 

katkit25:

k thank you and i have been here for a while @pooja195

pooja195:

My apologies, let me know when you have the answer!

katkit25:

ok

katkit25:

i got the answer it is 100

pooja195:

5*4=20/2=10 4*20=80 80+10=?

pooja195:

You're just slightly off

pooja195:

Did you get it?

katkit25:

i read my answer wrong srry it is 90

katkit25:

closeing this soon answer now or ask now before et gets closed just letting you guys know